# Current Pathology Foundation Models are unrobust to Medical Center Differences

### v1

@@ 2 Related Work
In simultaneous work from the TU Berlin BIFOLD group [25], batch effects in pathology foundation models were analyzed and shown. [26] analyzes rotation invariance. Tellez [21] quantified the effects of data augmentation and stain color normalization in pathology in the context of CNNs.
While finalizing this paper, a new relevant article discussing the measurement and optimization of robustness in pathology became available [27]
@@ 3.1 Robustness Index
  • $k$ is the number of nearest neighbors considered; in this work, a value of $k=50$ is used
  • $y_{j}$ is the biological class of the $j$-th nearest neighbor
  • $y_{i}$ is the biological class of the sample $i$
  • $c_{j}$ is the medical center of the $j$-th nearest neighbor
  • $c_{i}$ is the medical center of the sample $i$
  • $\mathbf{1}(\cdot)$ is the indicator function, which equals 1 if the condition is true and 0 otherwise
@@
As an example, consider an embedding spaced dominated by a confounder, say center, and only minutely influenced by the cancer type. Then, the denominator will be large, since most samples will be surrounded by other samples of the same center; the numerator is small since the embedding space is mostly organised by the confounder.
In an idealised scenario, the embedding space is blind to the confounding information and completely organised by the biological signal, yielding $R_{k}\rightarrow\infty$. In practice, it may not be feasible to completely remove center information, but one would like to see $R_{k}>>1$.
@@ 4.1 Classification Task: Tissue of Origin / Cancer Type
To build the dataset, for each available combination of center and cancer type, 10 WSIs are selected randomly. From each of the resulting WSIs, 10 informative foreground patches representing regular tissue were selected from a larger randomly generated set based on visual inspection, filtering out anomalies and low-information (white) patches. This resulted in a dataset of 2000 patches in total. In all experiments, 5-fold cross-validation is used to generate validation predictions for this whole dataset.
@@ 4.2 Control Classification Task: Medical Center Predictions
Table 1: Tissue Source Site (TSS), Short Name and Project Code Combinations
@@ 5.1 Embedding Space Structure and Robustness Index
Figure 1: Fraction of samples for which the k-th neighbor has the same cancer type (blue) or medical center (orange), in order of increasing robustness. See Appendix 12 for a complete overview of these results. For all models, closeness in the embedding space is strongly determined by whether the image comes from the same medical center. For all models except Virchow2, the medical center more strongly determines embedding proximity than the cancer type for the closest 200 neighbors.
@@ 5.2 Quantification of the Influence of Medical Center Differences on FM-based Prediction Performance
Figure 3: Fraction of same-center confounders for neighbors from the incorrectly predicted class (red); accuracy of tissue of origin / cancer type prediction (green); and accuracy of medical center prediction (blue). These figures are sorted (in natural reading order) by order of increasing center-robustness for selected FMs. All models show a substantial and significant influence of same-center confounders. See Appendix 9.3 for a complete overview of these results.
@@ 5.4 Relation between Prediction Performance and Robustness
Figure 7 shows the results for prediction of cancer type and medical center from embeddings. The top row shows prediction using knn with k=3 and logistic regression. For logistic regression (top right), we see that all models except SRA predict the medical center to a very high degree of accuracy: EXAONEPath, Phikon and Phikon-v2 predict the medical center with a cross-validated accuracy of 0.987, 0.987 and 0.993 respectively, the latter approaching perfect center prediction. See Table 2 for numerical results.
@@
For knn on the full embeddings (top left), the accuracy of both cancer type and medical center are reduced compared to logistic regression. There is a larger spread between the various results; and using knn, there is one model, Virchow2, that performs better on cancer type prediction than on medical center prediction, indicating a better relation between biologically relevant prediction performance and robustness.
The bottom two graphs show analogous results, but based on using 2D t-SNE coordinates as input rather than the full embeddings.
@@
5.5 Effect of Medical Center Influences on Logistic Regression

### v2

@@ 2 Related Work
In simultaneous work from the TU Berlin BIFOLD group [25], batch effects in pathology foundation models were analyzed and shown. [26] analyzes rotation invariance. Tellez [21] quantified the effects of data augmentation and stain color normalization in pathology in the context of CNNs, and [27] evaluated the effect of stain normalization in colorectal tissue classification.
[28] discusses the issue that hospitals are represented in WSI data, and looks into domain generalization for hospital-agnostic image representation learning; [29] also uses domain adaptation to reduce the influence of staining differences. [30] draws attention to the existence of medical center signatures in TCGA data specifically, and finds that deep neural networks can predict the acquisition site.
A main consequence of unrobustness is that predictive models based on unrobust representations can lead to biased predictions. [31] studies the impact of medical center signatures on deep learning model accuracy and bias. [18] notes that embeddings from SSL models tend to cluster by individual WSIs and proposes the EXAONEPath pathology FM to address this, which is included in our evaluation here.
While finalizing this paper, a new relevant article discussing the measurement and optimization of robustness in pathology became available [32].
@@ 3.1 Robustness Index
  • $k$ is the number of nearest neighbors considered; in this work, $k=50$
  • $y_{j}$ is the biological class of the $j$-th nearest neighbor; $y_{i}$ is the biological class of the sample $i$
  • $c_{j}$ is the medical center of the $j$-th nearest neighbor; $c_{i}$ is the medical center of the sample $i$
  • $\mathbf{1}(\cdot)$ is the indicator function: 1 if the condition is true, 0 otherwise
@@
As an example, consider an embedding spaced dominated by a confounder, say center, and only minutely influenced by the cancer type. Then, the denominator will be large, since most samples will be surrounded by other samples of the same center; the numerator is small since the embedding space is mostly organized by the confounder.
A robustness index of 1 means that biological and confounding information are represented equally strongly. In an idealized scenario, the embedding space is blind to the confounding information and completely organized by the biological signal, yielding $R_{k}=R_{max}$, where $R_{max}$ is determined by the random chance level for encountering a neighbor from the same medical center or confounding class. In practice, it may not be feasible to completely remove center information, but one would like to see $R_{k}>>1$. For the dataset used in this paper, the Robustness Index is expected to vary between 1/4 for random cancer type prediction in combination with perfect medical center prediction and 4 for the converse.
@@ 4.1 Classification Task: Tissue of Origin / Cancer Type
To build the dataset, for each available combination of center and cancer type, 10 WSIs are selected randomly. From each of the resulting WSIs, 10 informative foreground patches representing regular tissue were selected from a larger randomly generated set based on visual inspection, filtering out anomalies and low-information (white) patches. This resulted in a dataset of 2000 patches in total. We name this dataset TCGA-2k. In all experiments, 5-fold cross-validation is used to generate validation predictions for this whole dataset.
@@ 4.2 Control Classification Task: Medical Center Predictions
Table 1: Composition of the TCGA-2k dataset: Tissue Source Site (TSS), Short Name and Project Code Combinations
@@ 5.1 Embedding Space Structure and Robustness Index
Figure 1: Fraction of samples for which the k-th neighbor has the same cancer type (blue) or medical center (orange), in order of increasing robustness. See Appendix 12 for all results. For all models, closeness in embedding space is strongly determined by whether the image comes from the same medical center. For all models except Virchow2, the medical center more strongly determines embedding proximity than the cancer type for the nearest 200 neighbors.
@@ 5.2 Quantification of the Influence of Medical Center Differences on FM-based Prediction Performance
Figure 3: Fraction of same-center confounders for neighbors from the incorrectly predicted class (red); accuracy of tissue / cancer type prediction (green); accuracy of medical center prediction (blue), sorted by order of increasing center-robustness for selected FMs. All models show a substantial and significant influence of same-center confounders. See Appendix 9.3 for a complete overview.
@@ 5.4 Relation between Prediction Performance and Robustness
Figure 7 shows the results for prediction of cancer type and medical center from embeddings. The top row shows prediction using knn with k=3 and logistic regression. For logistic regression (top right), we see that all models except SRA predict the medical center to a very high degree of accuracy: EXAONEPath, Phikon and Phikon-v2 have cross-validated accuracies of 0.987, 0.987 and 0.993 respectively, the latter approaching perfect center prediction. See Table 2 for numerical results.
@@
For knn on the full embeddings (top left), the accuracy of cancer type and medical center are reduced compared to logistic regression. There is a larger spread between the various results; and using knn, there is one model, Virchow2, that performs better on cancer type prediction than on medical center prediction, indicating a better relation between biologically relevant prediction performance and robustness. The bottom two graphs show analogous results, but based on using 2D t-SNE coordinates as input rather than the full embeddings.
@@
5.5 Effect of Medical Center Influences on Regression